To state in interval notation the absolute value of 2x-2 is greater than or eq to 1/2 do I move the -2 and change the sign of the equality?
thank you!

Answer by stanbon(75887)

To state in interval notation the absolute value of 2x-2 is greater than or eq to 1/2 do I move the -2 and change the sign of the equality?
-------------
|2x-2|<=(1/2) means,
-1/2<=2x-2<=(1/2)
Add 2 along the line to get:
(3/2)<=2x<=5/2
Divide thru by 2 to get:
3/4<=x<=5/4
In interval notation this is written [3/4,5/4]
Cheers,
Stan H.